Consider the equation (x − 2)5€ = 0,(3) What is the convergence order of your Newtons iteration if the iteration converges? (4) Use the iteration formula in (2)
Answer #1

@ The ander of convergence of Meewfon-Raphson method is 2. 1 9 het f(x) = (x-2)5 ex . The Mewton-Raphson ifensation formula iTaking madulas both side we have Horsea | K+1 -2/2 |XK+2)(9–274 . (*x+3) - | ** 2) park-21 72x+3) ? : | Mart 2] = lock +21-*